OpenStudy (anonymous):

Simplify: x^2+2x/12x+54 - 3-x/8x+36

OpenStudy (anonymous):

x2+2x12x+54−3−x8x+36 =x(x+2)6(2x+9)−3−x4(2x+9) =2x(x+2)12(2x+9)−3(3−x)12(2x+9) =2x(x+2)−3(3−x)12(2x+9) =2x2+7x−912(2x+9) =(x−1)(2x+9)12(2x+9) =x−112

OpenStudy (anonymous):

http://inquus.org/updates/4e557b400b8b9ebaa88e66a4

OpenStudy (anonymous):

Thank youu!

OpenStudy (anonymous):

thanks u welcome canu help me... http://openstudy.com/study#/updates/5141f372e4b0f08cbdc993c2